A 1.00 mole sample of N2 gas at 0 ∘C is heated to 180 ∘C at constant pressure (1.00 atm).

1. Determine the change in internal energy.

2. Determine the work the gas does

3. Determine the heat added to it.
Answer #1

A)

Change in internal energy, dU = n(5/2)R dT

dU = 1 x (5/2) x 8.314 x 180 = 3741.3 J

B)

Work done, W = nRdT = 1 x 8.314 x 180 = 1496.52 J

C)

Heat, Q = dU + W = 5237.82 J

Comment in case any doubt please rate my answer ....